# 3cm3cm3cm3cm

### v1

@@ 4.2. From Berezinians to column determinants
$\overline{\phi}_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!),$
respectively.
@@
$\mathcal{C}=\left[\delta_{a,b}(z-w_{a})-\left(\sum_{i=1}^{m+n}\frac{x^{b}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}-\sum_{r=1}^{p+q}\frac{y^{a}_{r}\partial_{y^{b}_{r}}}{\partial_{z}-z_{r}}\right)\right]_{a,b=1,\ldots,d}.$
Lemma 4.3.
We have
@@
$\phi_{\bf z}\big{(}e_{ab}(z)\big{)}=\sum_{i=1}^{m+n}\frac{x^{a}_{i}\partial_{x^{b}_{i}}}{z-z_{p+q+i}}-\sum_{r=1}^{p+q}\frac{y^{b}_{r}\partial_{y^{a}_{r}}}{z-z_{r}}.$ (4.7)
We define
$F^{i}_{a,b}=\begin{cases}\,\,\delta_{a,b}(\partial_{z}-w_{a})&\mbox{if }\,i=0,\\ \displaystyle{\,\,\frac{y^{b}_{i}\partial_{y^{a}_{i}}}{z-z_{i}}}&\mbox{if }\,1\leq i\leq p+q,\\ \displaystyle{\,\frac{-x^{a}_{i-p-q}\partial_{x^{b}_{i-p-q}}}{z-z_{i}}}&\mbox{if }\,p+q+1\leq i\leq p+q+m+n,\end{cases}$
@@
$G^{i}_{a,b}=\begin{cases}\,\,\delta_{a,b}(z-w_{a})&\mbox{if }\,i=0,\\ \displaystyle{\,\,\frac{y^{a}_{i}\partial_{y^{b}_{i}}}{\partial_{z}-z_{i}}}&\mbox{if }\,1\leq i\leq p+q,\\ \displaystyle{\,\frac{-x^{b}_{i-p-q}\partial_{x^{a}_{i-p-q}}}{\partial_{z}-z_{i}}}&\mbox{if }\,p+q+1\leq i\leq p+q+m+n.\end{cases}$
We have
$\omega(F^{i}_{a,b})=G^{i}_{a,b}$ (4.8)
for all $0\leq i\leq{p+q+m+n}$ and $1\leq a,b\leq d$.
Note that ${\mathcal{L}}_{d}^{\bf w}$ is an amply invertible Manin matrix of type $(0^{d})$. Let $\pi\in\mathfrak{S}_{d}$ be the permutation given by $\pi(i)=d+1-i$ for $i=1,\ldots,d$. Clearly, $\pi^{2}=1$. By (4.7), Proposition 3.5 and Proposition 3.12, we obtain
@@
In view of (4.8), we deduce that $\omega\big{(}{\rm cdet}\,\phi_{\bf z}\,{\mathcal{L}}_{d}^{\bf w}\big{)}={\rm cdet}\,(\mathcal{C})$. ∎
Fix ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{p+m|q+n}$. For $u=z$ or $\partial_{z}$, let
$R_{p|q}(u)=\frac{(u-z_{1})\ldots(u-z_{p})}{(u-z_{p+1})\ldots(u-z_{p+q})}$
and
Proposition 4.4.
$(z-w_{1})\ldots(z-w_{d})\,\overline{\phi}_{\bf w}\,{\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)=\omega\left(\!R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})\cdot R_{p|q}(z)\!\right).$
Proof.
For $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$, we have
${\begin{array}[]{cll}\overline{\phi}_{\bf w}\big{(}E^{r}_{s}(z)\big{)}&=\sum_{a=1}^{d}\frac{(-1)^{|s|+1}\partial_{y^{a}_{r}}y^{a}_{s}}{z-w_{a}},\quad\overline{\phi}_{\bf w}\big{(}E^{r}_{p+q+j}(z)\big{)}&=\sum_{a=1}^{d}\frac{\partial_{y^{a}_{r}}\partial_{x^{a}_{j}}}{z-w_{a}},\\ \overline{\phi}_{\bf w}\big{(}E^{p+q+i}_{s}(z)\big{)}&=\sum_{a=1}^{d}\frac{(-1)^{|s|+1}x^{a}_{i}y^{a}_{s}}{z-w_{a}},\quad\overline{\phi}_{\bf w}\big{(}E^{p+q+i}_{p+q+j}(z)\big{)}&=\sum_{a=1}^{d}\frac{x^{a}_{i}\partial_{x^{a}_{j}}}{z-w_{a}}.\end{array}}$ (4.9)
@@
$D=\Big{[}\delta_{a,b}(z-w_{a})\Big{]}_{a,b=1,\ldots,d}\quad\text{and}\quad D^{\prime}=\Big{[}\delta_{i,j}(\partial_{z}-z_{i})\Big{]}_{i,j\in\mathbb{I}}.$
@@
$X:=\big{[}(-1)^{|i|}x^{a}_{i}\big{]}^{a=1,\ldots,d}_{i=1,\ldots,m+n}\quad\text{and}\quad P_{X}:=\big{[}\partial_{x^{a}_{i}}\big{]}^{a=1,\ldots,d}_{i=1,\ldots,m+n}$

### v2

@@ 4.1. Column determinants and Berezinians
$\varphi_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!),$
respectively. Furthermore, there is an anti-involution $\omega:\mathcal{D}\longrightarrow\mathcal{D}$ defined by
$x^{a}_{i}\mapsto\partial_{x^{a}_{i}},\quad\partial_{x^{a}_{i}}\mapsto x^{a}_{i},\quad y^{a}_{r}\mapsto\partial_{y^{a}_{r}},\quad\partial_{y^{a}_{r}}\mapsto y^{a}_{r},$ (4.13)
for $1\leq i\leq m+n$, $1\leq r\leq p+q$, and $1\leq a\leq d$. In light of Proposition 2.1, the map $\omega$ extends naturally to an anti-involution
For $u=z$ or $\partial_{z}$, let
$R_{p|q}(u)=\frac{(u-z_{1})\ldots(u-z_{p})}{(u-z_{p+1})\ldots(u-z_{p+q})}$
and
We have the following proposition. We defer its proof until Appendix A to avoid disrupting the flow of the paper.
Proposition 4.7.
Let ${\bf s}=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{p+m|q+n}$. Then
$(z-w_{1})\ldots(z-w_{d})\,\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)=\omega\left(\!R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})\cdot R_{p|q}(z)\!\right).$
The Cartan anti-involution $\varpi$ on ${\mathfrak{gl}}_{d}$, defined by (3.14), gives rise to an anti-involution $\varpi:U({\mathfrak{gl}}_{d}[t])\longrightarrow U({\mathfrak{gl}}_{d}[t])$, defined by
$\hskip 42.67912pt\varpi(A[k])=\varpi(A)[k]\qquad\mbox{for $A\in{\mathfrak{gl}}_{d}$ and $k\in{\mathbb{Z}}_{+}$}.$
Proposition 4.8 ([MTV1, Proposition 8.4]).
The anti-involution $\varpi:U({\mathfrak{gl}}_{d}[t])\longrightarrow U({\mathfrak{gl}}_{d}[t])$ fixes all elements of $\mathcal{B}_{d}^{\bf w}$.
Proposition 4.9.
For any ${\bm{b}}\in\mathcal{B}_{d}^{\bf w}$, $\omega\big(\phi_{\bf z}({\bm{b}})\big)=\phi_{\bf z}({\bm{b}})$.
Proof.
It is straightforward to verify that the diagram
[…]
commutes. This, together with Proposition 4.8, proves the proposition. ∎
The following theorem is called the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$. It not only relates the coefficients of $\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})$ to those of $\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$, but also gives an equality of algebras.
Theorem 4.10.
Let ${\bf s}=(0^{p},1^{q},0^{m},1^{n})\in\mathcal{S}_{p+m|q+n}$. We may write
$(z-w_{1})\ldots(z-w_{d})\ \varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$
Moreover, $\phi_{\bf z}(\mathcal{B}_{d}^{\bf w})=\varphi_{\bf w}\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)$.
Proof.
By Proposition 4.7,
$(z-w_{1})\ldots(z-w_{d})\,\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$
$\hskip 42.67912pt=R_{p|q}(\partial_{z})\cdot\omega\left(R_{p|q}(z)\cdot R^{\prime}_{m|n}(z)\cdot\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})\right)\cdot R_{p|q}(\partial_{z})^{-1}$
$\hskip 42.67912pt=R_{p|q}(\partial_{z})\left(\sum_{k=-\infty}^{p+m-q-n}\sum_{\ell=0}^{d}\omega\big(b_{k,\ell}^{{\bf w},{\bf z}}\big)z^{\ell}\partial_{z}^{k}\right)R_{p|q}(\partial_{z})^{-1}.$
By Proposition 4.9, $\omega\big(b_{k,\ell}^{{\bf w},{\bf z}}\big)=b_{k,\ell}^{{\bf w},{\bf z}}$ for all $k\leq p+m-q-n$ and $0\leq\ell\leq d$. This proves the first assertion. The last assertion clearly follows from the first one. ∎
Remark 4.11.
We have learned from an anonymous expert that the last assertion of Theorem 4.10 can be obtained directly from the results of [HM]. An outline of his approach is given as follows. Let $\pi$ be an automorphism on $\mathcal{D}$ defined by
$x^{a}_{i}\mapsto x^{a}_{i},\quad\partial_{x^{a}_{i}}\mapsto\partial_{x^{a}_{i}},\quad y^{a}_{r}\mapsto(-1)^{|r|+1}\partial_{y^{a}_{r}},\quad\partial_{y^{a}_{r}}\mapsto y^{a}_{r},$
for $1\leq i\leq m+n$, $1\leq r\leq p+q$, and $1\leq a\leq d$, and let $\vartheta$ be an automorphism on $U({\mathfrak{gl}}_{d}[t])$ defined by
$e_{ab}[k]\mapsto e_{ab}[k]-\delta_{a,b}\sum_{r=1}^{p+q}(-1)^{|r|}z_{r}^{k}$
for $1\leq a,b\leq d$ and $k\in{\mathbb{Z}}_{+}$. Note that $\vartheta$ restricts to an automorphism on $\mathcal{B}_{d}^{\bf w}$. As a consequence of [HM, Theorem 5.2],
$(\pi\circ\phi_{\bf z}\circ\vartheta)(\mathcal{B}_{d}^{\bf w})=(\pi\circ\varphi_{\bf w})({\mathcal{B}_{p+m|q+n}^{\bf z}}),$
which is equivalent to the last assertion of Theorem 4.10.
4.4. An application of the Bethe duality
The Weyl superalgebra $\mathcal{D}$ is a subalgebra of ${\rm End}(\mathcal{F})$. For ${\bf w}\in{\mathbb{C}}^{d}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$, Theorem 4.10 implies that the image of $\mathcal{B}_{d}^{\bf w}$ in ${\rm End}(\mathcal{F})$ under $\phi_{\bf z}$ coincides with the image of $\mathcal{B}_{p+m|q+n}^{\bf z}$} in ${\rm End}(\mathcal{F})$ under $\varphi_{\bf w}$. This makes it possible to study the action of $\mathcal{B}_{p+m|q+n}^{\bf z}$} on $\mathcal{F}$ via the action of $\mathcal{B}_{d}^{\bf w}$ on $\mathcal{F}$, and vice versa.
@@ 4.4. An application of the Bethe duality
${\bf d}_{\mathcal{F}}=\sum_{a=1}^{d}\left(\!\sum_{i=1}^{m}x_{i}^{a}\partial_{x^{a}_{i}}+\sum_{j=1}^{n}\eta_{j}^{a}\partial_{\eta^{a}_{j}}-\sum_{r=1}^{p}y_{r}^{a}\partial_{y^{a}_{r}}-\sum_{s=1}^{q}\zeta_{s}^{a}\partial_{\zeta^{a}_{s}}\!\right)\!.$
For any monomial $f\in\mathcal{F}$, we have ${\bf d}_{\mathcal{F}}(f)=kf$ for some $k\in{\mathbb{Z}}$. We define the degree of $f$ to be $k$. In particular, the degrees of $x^{a}_{i}$} and $\eta^{a}_{j}$} are $1$ while the degree of $y^{a}_{r}$ and $\zeta^{a}_{s}$ are 1-1. We call $\bf d$$_{\mathcal{F}}$ the degree operator on $\mathcal{F}$.
For $1\leq r\leq p$, $1\leq s\leq q$, $1\leq i\leq m$, and $1\leq j\leq n$, let $\mathcal{F}_{(r)}={\mathbb{C}}\big[y^{1}_{r},\ldots,y^{d}_{r}\big]$, $\mathcal{F}_{(p+s)}={\mathbb{C}}\big[\zeta^{1}_{s},\ldots,\zeta^{d}_{s}\big]$, $\mathcal{F}_{(p+q+i)}={\mathbb{C}}\big[x^{1}_{i},\ldots,x^{d}_{i}\big]$, and $\mathcal{F}_{(p+q+m+j)}={\mathbb{C}}\big[\eta^{1}_{j},\ldots,\eta^{d}_{j}\big]$. It is evident that as ${\mathfrak{gl}}_{d}$-modules,
$\mathcal{F}=\mathcal{F}_{(1)}\otimes\cdots\otimes\mathcal{F}_{(p+q+m+n)}.$ (4.14)
For $1\leq a\leq d$, let $\mathcal{F}^{(a)}={\mathbb{C}}\big[x^{a}_{i},\eta^{a}_{j},y^{a}_{r},\zeta^{a}_{s}\,\big|\,1\leq i\leq m,1\leq j\leq n,1\leq r\leq p,1\leq s\leq q\big]$. As ${\mathfrak{gl}}_{p+m|q+n}$-modules,
$\mathcal{F}=\mathcal{F}^{(1)}\otimes\cdots\otimes\mathcal{F}^{(d)}.$ (4.15)
For any weight $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ of a ${\mathfrak{gl}}_{p+m|q+n}$-module $M$, we denote by $M_{\mu}$ the $\mu$-weight space of $M$. The Cartan subalgebra ${\mathfrak{h}}_{{p+m|q+n}}$ of ${\mathfrak{gl}}_{p+m|q+n}$ acts on $\mathcal{F}$ as follows (see (4.10)): $E^{i}_{i}$ acts on $\mathcal{F}$ by $\sum_{a=1}^{d}y^{a}_{i}\partial_{y^{a}_{i}}-d,$ for $1\leq i\leq p$, and by $\sum_{a=1}^{d}\zeta^{a}_{i-p}\partial_{\zeta^{a}_{i-p}}+d,$ for $p+1\leq i\leq q$. Also, $E^{p+q+i}_{p+q+i}$ acts on $\mathcal{F}$ by $\sum_{a=1}^{d}x^{a}_{i}\partial_{x^{a}_{i}},$ for $1\leq i\leq m$, and by $\sum_{a=1}^{d}\eta^{a}_{i-m}\partial_{\eta^{a}_{i-m}},$ for $m+1\leq i\leq m+n$. Clearly, there is a weight space decomposition
$\mathcal{F}=\bigoplus_{\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}}\mathcal{F}_{\mu}.$ (4.16)
Suppose $p\not=0\not=m$. For $r\in{\mathbb{Z}}$, let
[…]
The space $V_{r}$ is an infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module (with respect to ${\bm{\sigma}}_{p}$), and the highest weight of $V_{r}$, appearing in (4.17), corresponds to a generalized partition of depth 1; see [CCL1, Section 2.5] and also [CLZ]. (Such a partition is called a generalized partition of length 1 in [CLZ].)
For $a=1,\ldots,d$, let $\gamma_{a}\in{\mathbb{Z}}$. If $\gamma_{a}\geq 0$, then we may identify $V_{\gamma_{a}}$ with the ${\mathfrak{gl}}_{p+m|q+n}$-submodule of the polynomial superalgebra $\mathcal{F}^{(a)}$ generated by the highest weight vector $(x^{a}_{1})^{\gamma_{a}}$. If $\gamma_{a}<0$, then we may identify $V_{\gamma_{a}}$ with the ${\mathfrak{gl}}_{p+m|q+n}$-submodule of $\mathcal{F}^{(a)}$ generated by the highest weight vector
In any case, $V_{\gamma_{a}}$ is the subspace of $\mathcal{F}^{(a)}$ spanned by all monomials of degree $\gamma_{a}$. Via (4.15), the tensor product $V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}}$ is a ${\mathfrak{gl}}_{p+m|q+n}$-submodule of $\mathcal{F}$. We thus obtain a direct sum decomposition of ${\mathfrak{gl}}_{p+m|q+n}$-modules:
$\mathcal{F}=\bigoplus_{\gamma_{1},\ldots,\gamma_{d}\in{\mathbb{Z}}}V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}}.$ (4.18)
Let ${\mathfrak{h}}_{d}$ denote the standard Cartan subalgebra of ${\mathfrak{gl}}_{d}$. The action of ${\mathfrak{h}}_{d}$ on $\mathcal{F}$ is determined by the action of $e_{aa}$ on $\mathcal{F}$, which is given by
$\sum_{i=1}^{m}x_{i}^{a}\partial_{x^{a}_{i}}+\sum_{j=1}^{n}\eta_{j}^{a}\partial_{\eta^{a}_{j}}-\sum_{r=1}^{p}y_{r}^{a}\partial_{y^{a}_{r}}-\sum_{s=1}^{q}\zeta_{s}^{a}\partial_{\zeta^{a}_{s}},$
for $1\leq a\leq d$ (see (4.9)). There is a weight space decomposition
$\mathcal{F}=\bigoplus_{\gamma\in{\mathfrak{h}}_{d}^{*}}\mathcal{F}_{\gamma}.$ (4.19)
For $1\leq i\leq p+q$ and $k\in-{\mathbb{Z}}_{+}$, let $\mathcal{F}_{(i)}^{k}$ be the subspace of $\mathcal{F}_{(i)}$ spanned by all monomials of degree $k$. For $1\leq j\leq m+n$ and $l\in{\mathbb{Z}}_{+}$, let $\mathcal{F}_{(p+q+j)}^{l}$ be the subspace of $\mathcal{F}_{(p+q+j)}$ spanned by all monomials of degree $l$. Evidently, $\mathcal{F}_{(i)}^{k}$ and $\mathcal{F}_{(p+q+j)}^{l}$ are finite-dimensional ${\mathfrak{gl}}_{d}$-modules. We have the following direct sum decompositions of ${\mathfrak{gl}}_{d}$-modules:
$\mathcal{F}_{(i)}=\bigoplus_{k\in-{\mathbb{Z}}_{+}}\mathcal{F}_{(i)}^{k}\qquad\text{and}\qquad\mathcal{F}_{(p+q+j)}=\bigoplus_{l\in{\mathbb{Z}}_{+}}\mathcal{F}_{(p+q+j)}^{l}.$
It is well known that $\mathcal{F}_{(i)}^{k}$ and $\mathcal{F}_{(p+q+j)}^{l}$ are irreducible ${\mathfrak{gl}}_{d}$-modules for $k\in-{\mathbb{Z}}_{+}$ and $l\in{\mathbb{Z}}_{+}$ (see, for instance, [CLZ, Theorem 3.3] by setting exactly one of $p,q,m,n$ (given there) to 1 and the others to 0). By (4.14), $\mathcal{F}$ decomposes into a direct sum of ${\mathfrak{gl}}_{d}$-modules
$\mathcal{F}=\bigoplus_{\begin{subarray}{c}k_{1},\ldots,k_{p+q}\in-{\mathbb{Z}}_{+},\\ k_{p+q+1},\ldots,k_{p+q+m+n}\in{\mathbb{Z}}_{+}\end{subarray}}\mathcal{F}_{(1)}^{k_{1}}\otimes\cdots\otimes\mathcal{F}_{({p+q+m+n})}^{k_{p+q+m+n}}.$ (4.20)
Proposition 4.12.
For $\gamma_{1},\ldots,\gamma_{d}\in{\mathbb{Z}}$, let $\gamma=\sum_{a=1}^{d}\gamma_{a}\epsilon_{a}\in{\mathfrak{h}}_{d}^{*}$. Let $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ be any weight of $V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}}$. Then
$(V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}})_{\mu}=\big(\mathcal{F}_{(1)}^{\overline{\mu}_{1}}\otimes\cdots\otimes\mathcal{F}_{({p+q+m+n})}^{\overline{\mu}_{p+q+m+n}}\big)_{\gamma},$
where
$\overline{\mu}_{i}=\begin{cases}\,\mu(E^{i}_{i})+(-1)^{|i|}d&\mbox{if}\ \ 1\leq i\leq p+q;\\ \,\mu(E^{i}_{i})\ \ &\mbox{if}\ \ p+q+1\leq i\leq p+q+m+n.\end{cases}$ (4.21)
Proof.
By (4.18) and (4.19), we have
$\mathcal{F}_{\gamma}=V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}}.$
Also, (4.16) and (4.20) yield
$\mathcal{F}_{\mu}=\mathcal{F}_{(1)}^{\overline{\mu}_{1}}\otimes\cdots\otimes\mathcal{F}_{({p+q+m+n})}^{\overline{\mu}_{p+q+m+n}}.$
Consequently,
$(V_{\gamma_{1}}\otimes\cdots\otimes V_{\gamma_{d}})_{\mu}=\mathcal{F}_{\gamma}\cap\mathcal{F}_{\mu}=\big(\mathcal{F}_{(1)}^{\overline{\mu}_{1}}\otimes\cdots\otimes\mathcal{F}_{\overline{\mu}_{p+q+m+n}}^{(p+q+m+n)}\big)_{\gamma},$
as asserted. ∎
For ${\mathfrak{gl}}_{p+m|q+n}$-modules $M_{1},\ldots,M_{d}$ and ${\bf w}=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}$, we write
Theorem 4.13.
Suppose $p\not=0\not=m$. For $a=1,\ldots,d$, let $L_{a}=V_{\gamma_{a}}$ for some $\gamma_{a}\in{\mathbb{Z}}$. For any weight $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ of $\underline{L}$, ${\bf z}\in{\bm{X}}_{\!p+q+m+n}$, and ${\bf w}\in{\bm{X}}_{\!d}$, we have:
  1. (i)
    $\underline{L}({\bf w})_{\mu}$ is a cyclic $\mathcal{B}_{p+m|q+n}^{\bf z}$}-module.
  2. (ii)
    $\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{\underline{L}({\bf w})_{\mu}}$ is a Frobenius algebra.
  3. (iii)
    $\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{\underline{L}({\bf w})_{\mu}}$ is a maximal commutative subalgebra of ${\rm End}(\underline{L}({\bf w})_{\mu})$ of dimension $\dim\!\left(\underline{L}({\bf w})_{\mu}\right)$.
  4. (iv)
    Every eigenspace of $\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{\underline{L}({\bf w})_{\mu}}$ is one-dimensional.
Moreover, $\mathcal{B}_{p+m|q+n}^{\bf z}$} is diagonalizable with a simple spectrum on $\underline{L}({\bf w})_{\mu}$ for generic $\bf z$} and generic $\bf w$}.
Proof.
Let $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ be a weight of $\underline{L}$. For $i\in\mathbb{I}$, let $N_{i}=\mathcal{F}_{(i)}^{\overline{\mu}_{i}}$, where $\overline{\mu}_{i}$ is defined as in (4.21). Write ${\bf z}=(z_{1},\ldots,z_{p+q+m+n})$. We set $\underline{N\mkern-4.0mu}\mkern 4.0mu=N_{1}\otimes\cdots\otimes N_{p+q+m+n}$ and $\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})=N_{1}(z_{1})\otimes\cdots\otimes N_{p+q+m+n}(z_{p+q+m+n})$. In view of Theorem 4.10 and Proposition 4.12,
$\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{\underline{L}({\bf w})_{\mu}}=(\mathcal{B}_{d}^{\bf w})_{\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}}.$ (4.22)
By Theorem 3.3, $\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}$ is a cyclic $\mathcal{B}_{d}^{\bf w}$-module as $\mathcal{B}_{d}^{\bf w}$ preserves weight spaces, and by Theorem 3.14, $(\mathcal{B}_{d}^{\bf w})_{\underline{N\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}}$ is a Frobenius algebra. Hence, (i) and (ii) follow from (4.22). Also, (iii) and (iv) are immediate consequences of [Lu, Lemma 1.3] (cf. Remark 3.16). The last statement follows from (4.22) and Theorem 3.4. ∎
Remark 4.14.
Theorem 4.13 suggests that the super analog of Conjecture 3.13 should be true for the general linear Lie superalgebra.
Conjecture 4.15.
Suppose $p$, $q$, $m$, and $n$ are not all zero. For $a=1,\ldots,d$, let $L_{a}$ be a unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module (with respect to ${\bm{\sigma}}_{p}$). For any weight $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ of $\underline{L}$, ${\bf z}\in{\bm{X}}_{\!p+q+m+n}$, and ${\bf w}\in{\bm{X}}_{\!d}$, parts (i)–(iv) of Theorem 4.13 hold for the Bethe algebra $({\mathcal{B}_{p+m|q+n}^{\bf z}})_{\underline{L}({\bf w})_{\mu}}$. Moreover, $\mathcal{B}_{p+m|q+n}^{\bf z}$} is diagonalizable with a simple spectrum on $\underline{L}({\bf w})_{\mu}$ for generic $\bf z$} and generic $\bf w$}.
Remark 4.16.
An analog of Conjecture 4.15 when the entries of $\bf z$} are all zero is studied in [CCL2].
Suppose $p\not=0\not=m$. For $r\in{\mathbb{Z}}$, let
$W_{r}=\begin{cases}L_{p+m|0}(r\epsilon_{p+1}-\mathbbm{1}_{p})&\quad\mbox{if $r\geq 0$,}\\ L_{p+m|0}(r\epsilon_{p}-\mathbbm{1}_{p})&\quad\mbox{if $r<0$}.\end{cases}$
The following provides evidence for Conjecture 3.13.
Corollary 4.17.
For $a=1,\ldots,d$, let $L_{a}=W_{\gamma_{a}}$ for some $\gamma_{a}\in{\mathbb{Z}}$. For any weight $\mu\in{\mathfrak{h}}_{p+m}^{*}$ of $\underline{L}$, ${\bf z}\in{\bm{X}}_{\!p+m}$, and ${\bf w}\in{\bm{X}}_{\!d}$, parts (i)–(iv) of Theorem 4.13 hold for the Bethe algebra $(\mathcal{B}_{p+m}^{\bf z})_{\underline{L}({\bf w})_{\mu}}$. Moreover, $\mathcal{B}_{p+m}^{\bf z}$ is diagonalizable with a simple spectrum on $\underline{L}({\bf w})_{\mu}$ for generic $\bf z$} and generic $\bf w$}.
Proof.
This is an immediate consequence of Theorem 4.13 by taking $q=n=0$. ∎
Acknowledgments. The first author was partially supported by NSTC grants 113-2115-M-006-010 and 114-2115-M-006-005. The second author was partially supported by NSTC grant 112-2115-M-006-015-MY2. The authors are grateful to an anonymous expert for valuable comments on Theorem 4.10 (see Remark 4.11).
Appendix A The proof of Proposition 4.7
The appendix is devoted to proving Proposition 4.7. We keep the notations defined in Section 4. Let
$\mathcal{C}=\left[\delta_{a,b}(z-w_{a})+\sum_{r=1}^{p+q}\frac{y^{a}_{r}\partial_{y^{b}_{r}}}{\partial_{z}-z_{r}}-\sum_{i=1}^{m+n}\frac{x^{b}_{i}\partial_{x^{a}_{i}}}{\partial_{z}-z_{p+q+i}}\right]_{a,b=1,\ldots,d}.$
Lemma A.1.
@@
$\phi_{\bf z}\big(e_{ab}(z)\big)=-\sum_{r=1}^{p+q}\frac{y^{b}_{r}\partial_{y^{a}_{r}}}{z-z_{r}}+\sum_{i=1}^{m+n}\frac{x^{a}_{i}\partial_{x^{b}_{i}}}{z-z_{p+q+i}}.$ (A.1)
For $i\in\mathbb{I}\cup\{0\}$ and $1\leq a,b\leq d$, we define
$F^{i}_{a,b}=\begin{cases}\,\,\delta_{a,b}(\partial_{z}-w_{a})&\mbox{if }\,i=0,\\ \displaystyle{\,\,\frac{y^{b}_{i}\partial_{y^{a}_{i}}}{z-z_{i}}}&\mbox{if }\,1\leq i\leq p+q,\\ \displaystyle{\,\frac{-x^{a}_{i-p-q}\partial_{x^{b}_{i-p-q}}}{z-z_{i}}}&\mbox{otherwise},\end{cases}$
@@
$G^{i}_{a,b}=\begin{cases}\,\,\delta_{a,b}(z-w_{a})&\mbox{if }\,i=0,\\ \displaystyle{\,\,\frac{y^{a}_{i}\partial_{y^{b}_{i}}}{\partial_{z}-z_{i}}}&\mbox{if }\,1\leq i\leq p+q,\\ \displaystyle{\,\frac{-x^{b}_{i-p-q}\partial_{x^{a}_{i-p-q}}}{\partial_{z}-z_{i}}}&\mbox{otherwise}.\end{cases}$
We see that
$\omega(F^{i}_{a,b})=G^{i}_{a,b}.$ (A.2)
Note that ${\mathcal{L}}_{d}^{\bf w}$ is an amply invertible Manin matrix of type $(0^{d})$. Let $\pi\in\mathfrak{S}_{d}$ be the permutation given by $\pi(i)=d+1-i$ for $i=1,\ldots,d$. Clearly, $\pi^{2}=1$. By (A.1), Proposition 4.5 and Proposition 4.6, we obtain
@@
In view of (A.2), we deduce that $\omega\big(\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})\big)={\rm cdet}(\mathcal{C})$. ∎
Now we are ready to prove Proposition 4.7. For $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$, we have
$\begin{array}[]{cll}\varphi_{\bf w}\big(E^{r}_{s}(z)\big)\hskip-28.45274pt&=\sum_{a=1}^{d}\frac{(-1)^{|s|+1}\partial_{y^{a}_{r}}y^{a}_{s}}{z-w_{a}},\quad\varphi_{\bf w}\big(E^{r}_{p+q+j}(z)\big)\hskip-8.5359pt&=\sum_{a=1}^{d}\frac{\partial_{y^{a}_{r}}\partial_{x^{a}_{j}}}{z-w_{a}},\\ \varphi_{\bf w}\big(E^{p+q+i}_{s}(z)\big)\hskip-8.5359pt&=\sum_{a=1}^{d}\frac{(-1)^{|s|+1}x^{a}_{i}y^{a}_{s}}{z-w_{a}},\quad\varphi_{\bf w}\big(E^{p+q+i}_{p+q+j}(z)\big)\hskip-8.5359pt&=\sum_{a=1}^{d}\frac{x^{a}_{i}\partial_{x^{a}_{j}}}{z-w_{a}}.\end{array}$} (A.3)
@@
$D=\Big[\delta_{a,b}(z-w_{a})\Big]_{a,b=1,\ldots,d}\qquad\text{and}\qquad D^{\prime}=\Big[\delta_{i,j}(\partial_{z}-z_{i})\Big]_{i,j\in\mathbb{I}}.$
@@
$X:=\big[(-1)^{|p+q+i|}x^{a}_{i}\big]^{a=1,\ldots,d}_{i=1,\ldots,m+n}\qquad\text{and}\qquad P_{X}:=\big[\partial_{x^{a}_{i}}\big]^{a=1,\ldots,d}_{i=1,\ldots,m+n}$